3. What is the mass in grams of 6.022 x 1023 molecules of CO2? A. 22 g B. 44 g C. 668 D. 28 g
Answer #1

We have

molecules of CO2 = 6.022 \times 1023 molecules

We know that

1 mole of substance contains Avogadros number of molecules

Avogadros number = 6.022 \times 1023

that is

1 mole = 6.022 \times 1023 molecules

then Moles of CO2 = given number of molecules / Avogadros number

= 6.022 \times 1023 molecules /6.022 \times 1023 = 1 mole

Now we have 1 mole of CO2

lets calculate its mass by calculating its atoms atomic masses

then mass of CO2 = 1 \times 12 g/mol + 2 \times 16 g/mol

= 44 g

So answer is 44 g ....Option B is correct